OpenStudy (anonymous):

Please help! Solve the equation: 9 = -7m + 1 - 6 Please show all of the steps on how to solve this so I will know how to do it. :)

OpenStudy (ivettef365):

first you can start by combining the 1 - 6 so you have 9 = -7m -5 (1-6=-5) then add 5 to both sides as follows: 9+5 = -7m -5+5 (now on this side -5+5 cancels) and you have 14 = -7m

OpenStudy (ivettef365):

now divide by -7 on both sides as follows: 14 = -7m ---- ---- -7 -7

OpenStudy (ivettef365):

can you take it from there ?

OpenStudy (anonymous):

nice explaining ivette

OpenStudy (anonymous):

I got it! Thanks a lot for the help! :)

OpenStudy (ivettef365):

yw :)

OpenStudy (anonymous):

-14=7